\begin{thebibliography}{999}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\newcommand{\bb}{\bibitem}

{ 
\bibitem{Jeong:2011as}
Jeong, D.; Schmidt, F.; Hirata, C.M.
{Large-scale clustering of galaxies in general relativity.}
\emph{Phys. Rev. D} \textbf{2012}, {\emph{85}}, 023504.
%MDPI: We restored the title of all references
%doi:10.1103/PhysRevD.85.023504
%[arXiv:1107.5427 [astro-ph.CO]].
%253 citations counted in INSPIRE as of 29 Jun 2023

\bibitem{Schmidt:2012ne}
Schmidt, F.; Jeong, D.
Cosmic Rulers.
\emph{Phys. Rev. D} \textbf{2012}, \textit{86}, 083527.
%doi:10.1103/PhysRevD.86.083527
%[arXiv:1204.3625 [astro-ph.CO]].
%103 citations counted in INSPIRE as of 29 Jun 2023

\bibitem{Kehagias:2013yd}
Kehagias, A.; Riotto, A.
Symmetries and Consistency Relations in the Large Scale Structure of the Universe.
\emph{Nucl. Phys. B} \textbf{2013}, \textit{873}, 514--529.
%doi:10.1016/j.nuclphysb.2013.05.009
%[arXiv:1302.0130 [astro-ph.CO]].
%151 citations counted in INSPIRE as of 29 Jun 2023

\bibitem{Bertacca:2014dra}
Bertacca, D.; Maartens, R.; Clarkson, C.
Observed galaxy number counts on the lightcone up to second order: I. Main result.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2014},  \textit{9}, 037.
%doi:10.1088/1475-7516/2014/09/037
%[arXiv:1405.4403 [astro-ph.CO]].
%86 citations counted in INSPIRE as of 29 Jun 2023

\bibitem{Kehagias:2015tda}
Kehagias, A.; Dizgah, A.M.; Na, J.N.; Perrier, H.; Riotto, A.
A Consistency Relation for the Observed Galaxy Bispectrum and the Local non-Gaussianity from Relativistic Corrections.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2015}, \textit{8}, 18.
%doi:10.1088/1475-7516/2015/08/018
%[arXiv:1503.04467 [astro-ph.CO]].
%43 citations counted in INSPIRE as of 29 Jun 2023

\bibitem{Ginat:2021nww}
Ginat, Y.B.; Desjacques, V.; Jeong, D.; Schmidt, F.
Covariant decomposition of the non-linear galaxy number counts and their monopole.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2021}, \textit{12},  31.
%doi:10.1088/1475-7516/2021/12/031
%[arXiv:2106.15139 [astro-ph.CO]].
%4 citations counted in INSPIRE as of 29 Jun 2023
}


\bibitem{Yoo:2009au}
Yoo, J.; Fitzpatrick, A.L.; Zaldarriaga, M.
Three-point correlation of the Lyman-alpha forest: An optimal redshift space distortion estimator.
\emph{Phys. Rev. D} \textbf{2009}, \textit{80}, 083514.
%arXiv:0907.0707 [astro-ph.CO].

\bibitem{Yoo:2010ni}
Yoo, J.
A new relativistic N-body code for the clustering of cosmic neutrinos.
\emph{Phys. Rev. D} \textbf{2010}, \textit{82}, 083508.
%arXiv:1009.3021 [astro-ph.CO].

\bibitem{Challinor:2011bk}
Challinor, A.; Lewis, A.
The linear power spectrum of observed source number counts.
\emph{Phys. Rev. D} \textbf{2011}, \textit{84}, 043516.
%arXiv:1105.5292 [astro-ph.CO].

\bibitem{Bonvin:2011bg}
Bonvin, C.; Durrer, R.
What galaxy surveys really measure.
\emph{Phys. Rev. D } \textbf{2011}, \textit{84}, 063505.
%doi:10.1103/PhysRevD.84.063505.
%[arXiv:1105.5280 [astro-ph.CO]].
%384 citations counted in INSPIRE as of 09 Jun 2023

\bibitem{Grimm:2020bmv}
Grimm, N.; Scaccabarozzi, F.; Yoo, J.; Biern, S.G.; Gong, J.-O.
Precision cosmology with overlapping surveys: The importance of volume and cross-correlations.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2020}, \textit{11}, 64.
%arXiv:2005.06484 [astro-ph.CO].

\bibitem{Scaccabarozzi:2018iyz}
Scaccabarozzi, F.; Yoo, J.; Biern, S.G.
Cross-correlation of future weak lensing surveys and Planck lensing data.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2018}, \textit{10}, 24.
%arXiv:1807.09796 [astro-ph.CO].

\bibitem{Castorina:2021ihl}
Castorina, E.; Dio, E.D.
Observing the cosmic acceleration with the Kilo-Degree Survey.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2022}, \textit{1}, 61.
%arXiv:2106.08857 [astro-ph.CO].

\bibitem{1}
Fanizza, G.; Gasperini, M.; Marozzi, G.; Veneziano, G.
Generalized covariant prescriptions for averaging cosmological observables.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2020}, {\em 2}, 017.
%doi:10.1088/1475-7516/2020/02/017
%[arXiv:1911.09469 [gr-qc]].

\bibitem{2}
Ellis, G.
Relativistic cosmology.
\emph{Gen. Rel. Grav.}  \textbf{2009}, {\em 41}, 581--660.
%doi:10.1007/s10714-009-0760-7
%207 citations counted in INSPIRE as of 20 Feb 2020

\bibitem{3}
Ellis, G.; Nell, S.D.; Maartens, R.; Stoeger, W.R.; Whitman, A.P. Ideal observational cosmology.
\emph{Phys. Rep.} {\bf 1985}, \emph{124}, 315--417.

\bibitem{4}{Fleury, P.; Clarkson, C.; Maartens, R.} %MDPI: Please add the journal title
How does the cosmic large-scale structure bias the Hubble diagram?  
 \emph{J. Cosmol. Astropart. Phys.} \textbf{2017}, {\em 1703}, 062.
  %%CITATION = doi:10.1088/1475-7516/2017/03/062;%%
  %30 citations counted in INSPIRE as of 21 Nov 2019

\bibitem{5}{Dio, E.D.; Durrer, R.; Marozzi, G.; Montanari, F.} Galaxy number counts to second order and their bispectrum. \emph{J. Cosmol. Astropart. Phys.} {\textbf{2014}}, {\emph{1412}}, 17;
   Erratum in \emph{J. Cosmol. Astropart. Phys.} \textbf{2015}, {\emph{1506}}, E01.
   %MDPI: Please add the journal title of refs. 18,20-22,24 and 25.
   
     %%CITATION = doi:10.1088/1475-7516/2014/12/017, 10.1088/1475-7516/2015/06/E01;%%
  %66 citations counted in INSPIRE as of 21 Nov 2019
  
\bibitem{6}Gasperini, M.; Marozzi, G.; Nugier, F.; Veneziano, G. Light-cone averaging in cosmology: Formalism and applications. \emph{J. Cosmol. Astropart. Phys.} 
\textbf{2011}, {\emph{1107}},  008.
  %%CITATION = doi:10.1088/1475-7516/2011/07/008;%%
  %87 citations counted in INSPIRE as of 21 Nov 2019

\bibitem{7}{Ben-Dayan, I.; Gasperini, M.; Marozzi, G.; Nugier, F.; Veneziano, G.} Backreaction on the luminosity-redshift relation from gauge invariant light-cone averaging. %MDPI: Please add title.
 \emph{J. Cosmol. Astropart. Phys.} {\textbf{2012}}, {\emph{1204}},  36.
  %%CITATION = doi:10.1088/1475-7516/2012/04/036;%%
  %60 citations counted in INSPIRE as of 21 Nov 2019
  
\bibitem{8}{Fleury, P.; Nugier, F.; Fanizza, G.} Geodesic-light-cone coordinates and the Bianchi I spacetime. \emph{J. Cosmol. Astropart. Phys.} \textbf{2016}, {\em 6},  008.
  %%CITATION = doi:10.1088/1475-7516/2016/06/008;%%
  %16 citations counted in INSPIRE as of 21 Nov 2019

\bibitem{9}{Ben-Dayan, I.; Gasperini, M.; Marozzi, G.; Nugier, F.; Veneziano, G.} Average and dispersion of the luminosity-redshift relation in the concordance model. \emph{J. Cosmol. Astropart. Phys.} \textbf{2013}, {\em 1306},  2.
  %%CITATION = doi:10.1088/1475-7516/2013/06/002;%%
  %91 citations counted in INSPIRE as of 21 Nov 2019

\bibitem{10}
Fanizza, G.; Gasperini, M.; Marozzi, G.; Veneziano, G.
An exact Jacobi map in the  geodesic light-cone gauge.
\emph{J. Cosmol. Astropart. Phys.} \textbf{2013}, {\em 11}, 019.
%doi:10.1088/1475-7516/2013/11/019
%[arXiv:1308.4935 [astro-ph.CO]].
%71 citations counted in INSPIRE as of 20 Feb 2020

\bibitem{11}{Fanizza, G.; Gasperini, M.; Marozzi, G.; Veneziano, G.} Observation angles, Fermi coordinates, and the Geodesic-Light-Cone gauge. 
\emph{J. Cosmol. Astropart. Phys.} \textbf{2019}, {\em 1}, 4.

\bibitem{12}{Mitsou, E.; Scaccabarozzi, F.; Fanizza, G.} Observed Angles and Geodesic Light-Cone Coordinates. \emph{Class. Quantum Grav.} \textbf{2018}, {\em 35}, 107002.

\bibitem{Fonseca:2023uay}
Fonseca, J.; Zazzera, S.; Baker, T.; Clarkson, C. The observed number counts in luminosity distance space. \emph{arXiv} \textbf{2023},
arXiv:2304.14253.
%0 citations counted in INSPIRE as of 09 Jun 2023


\end{thebibliography}

\documentclass[preprints,communication,accept,pdftex,moreauthors]{Definitions/mdpi} 
\firstpage{1} 
\pubvolume{1}
\issuenum{1}
\articlenumber{0}
\pubyear{2023}
\copyrightyear{2023}
\externaleditor{Academic Editor: Firstname Lastname}
\datereceived{13 June 2023} 
\daterevised{3 July 2023} % Comment out if no revised date
\dateaccepted{7 July 2023} 
\datepublished{ } 
\hreflink{https://doi.org/} % If needed use \linebreak
\pdfoutput=1 % Uncommented for upload to arXiv.org

\def\laq{~\raise 0.4ex\hbox{$<$}\kern -0.8em\lower 0.62ex\hbox{$\sim$}~}
\def\gaq{~\raise 0.4ex\hbox{$>$}\kern -0.7em\lower 0.62ex\hbox{$\sim$}~}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def \ga {\gamma}

\def\beq{\begin{equation}}
\def\eeq{\end{equation}}
\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}

\def \pa {\partial}
\def \ra {\rightarrow}
\def \ti {\widetilde}

\newcommand{\Mcal}{\mathcal M}

\def\laq{~\raise 0.4ex\hbox{$<$}\kern -0.8em\lower 0.62ex\hbox{$\sim$}~}
\def\gaq{~\raise 0.4ex\hbox{$>$}\kern -0.7em\lower 0.62ex\hbox{$\sim$}~}

\def\beq{\begin{equation}}
\def\eeq{\end{equation}}
\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}
\def\bean{\begin{eqnarray*}}
\def\eean{\end{eqnarray*}}

\def \pa {\partial}
\def \ra {\rightarrow}
\def \ti {\widetilde}

\def \Da {\Delta}
\def \da {\delta}

\def \a {\alpha}

\def \ga {\gamma}

\def \da {\delta}

\def \Om {\Omega}

\def \U{\Upsilon}

\Title{A Simple, Exact Formulation of Number Counts in the Geodesic-Light-Cone Gauge}

\TitleCitation{Number Counts in the GLC Gauge}
\Author{{Giuseppe} %MDPI: Please carefully check the accuracy of names and affiliations. 
 Fanizza $^{1,\dagger}$, Maurizio Gasperini $^{2,3,}$*$^{,\dagger}$ and Giovanni Marozzi $^{4,5,\dagger}$}

\AuthorNames{Giuseppe Fanizza, Maurizio Gasperini and Giovanni Marozzi}

\AuthorCitation{{Fanizza,} %MDPI: Please carefully check the accuracy of names and affiliations. 
 G.; Gasperini, M.; Marozzi, G.}
\address{%
$^{1}$ \quad Instituto de Astrofis\'ica e Ci\^encias do Espa\c{c}o,
Faculdade de Ci\^encias da Universidade de Lisboa,
Edificio C8, Campo Grande, {P-1740-016} Lisbon, Portugal; gfanizza@fc.ul.pt\\
$^{2}$ \quad Dipartimento di Fisica, Universit\`a di Bari, 
Via G. Amendola 173, 70126 Bari, Italy\\
$^{3}$ \quad Istituto Nazionale di Fisica Nucleare, {Sezione di Bari}, 70126 Bari, Italy\\
$^{4}$ \quad Dipartimento di Fisica, Universit\`a di Pisa, Largo B. Pontecorvo 3, 56127 Pisa, Italy; giovanni.marozzi@unipi.it\\
$^{5}$ \quad Istituto Nazionale di Fisica Nucleare, {Sezione di Pisa}, 56127 Pisa, Italy}

\corres{Correspondence: gasperini@ba.infn.it}

\firstnote{These authors contributed equally to this work.} 
\abstract{In this article, we compare different formulations of the number count prescription using the convenient formalism of the Geodesic-Light-Cone gauge. We then find a simple, exact, and very general expression of such a prescription which is suitable for generalised applications.}

\keyword{\textls[-15]{{observational} cosmology; covariant averages of astrophysical variables; light cone gauge}
\\ \\ 
{\bf Published in} the Special Issue on {\em ``Universe: Feature Papers 2023 -- Cosmology"}, edited by K. Bamba ({\em Universe} {\bf 9}, 327 (2023)).
\\ \\ 
{\bf Preprint}: BA-TH/808-23  
\\ \\ 
{\bf Availabe at}: https://www.mdpi.com/2218-1997/9/7/327}
\begin{document}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\newpage

%%%%%%%%%%%%%%%%%%%%%%%%%
% INIZIO TESTO LAVORO
%%%%%%%%%%%%%%%%%%%%%%%%%
%MDPI: Please confirm whether you need to add section titles
{Galaxy} number counts represent a very useful tool for understanding the Large-Scale Structure (LSS) of our universe and testing the underlying cosmological models. Indeed, the information provided by such a tool on the distribution and overall properties of galaxies can be interpreted as a probe for the associated distribution of dark matter, and thereby as a test of the early cosmological dynamics based on late time observations {({see} \cite{Jeong:2011as,Schmidt:2012ne,Kehagias:2013yd,Bertacca:2014dra,Kehagias:2015tda,Ginat:2021nww} and references therein)}.
%MDPI: We removed the bold. Please confirm this revision.

One of the first studies involving galaxy number count was presented in a paper discussing the angular auto-correlation of the LSS \cite{Yoo:2009au} and its cross-correlation with the  anisotropies of the Cosmic Microwave Background (CMB) temperature. Later, this was generalized to include linear relativistic corrections \cite{Yoo:2010ni,Bonvin:2011bg,Challinor:2011bk}.

More recently, further developments have been achieved. Among the  most relevant, we should mention the computation of the theoretically expected galaxy power spectrum \cite{Grimm:2020bmv} based on the expression of galaxy number counts containing all relativistic corrections (i.e., at the source and observer positions and along the light cone). This advance complements the evaluation of the galaxy two-point correlation function presented in \cite{Scaccabarozzi:2018iyz}. In addition, \cite{Castorina:2021ihl} computed the ``observationally expected'' galaxy power spectrum by taking into account effects such as lensing magnification which are important in principle for the evaluation of the power spectrum multipoles, with the results showing that this approach can correct the leading distortion terms due to the redshift at most at the ten percent level.

On the other hand, in a previous paper \cite{1} we presented a new class of covariant prescriptions for averaging astrophysical variables on spatial regions typical of the chosen sources and intersecting the past light-cone of a given observer. In this short communication, we show that with the appropriate choice of ingredients, our integral prescription can exactly reproduce the number count integral introduced long ago as a useful tool in the general context of observational cosmology (see, e.g., \cite{2,3}).
In addition, it can reproduce the average integrals recently used in \cite{4,5} as essential ingredients to obtain a reliable (i.e., observationally compatible) prescription for galaxy number counts and for their above-mentioned applications.

Let us start by considering the (possibly realistic) experimental situation in which the sources are not exactly confined on a given space such as a hypersurface (defined for instance by a scalar field $B(x)$ such that $B(x)=B_s=${const}), and are instead localized inside an (arbitrarily) extended space-time layer corresponding to the interval  $B_s\le B\le B_s+\Delta B_s$. In thais case, the average is characterized by the following covariant integral prescription~\cite{1}:
\beq
I(\Da B_s)= \int_{\Mcal_4}d^4x\,
\sqrt{-g}\, \rho\,\da (V_o-V) \Theta(B_s+ \Da B_s-B) \Theta(B-B_s)\,{\pa^\mu A \pa_\mu V\over
\left|\pa_\a A \pa^\a A \right|^{1/2}}\,.
\label{1}
\eeq
{Here,} %MDPI: Please confirm whether you need to add indentation. The following highlights are the same.
 $\rho(x)$ is a scalar field that specifies an appropriate physical weight factor associated with the averaged sources, $V(x)$ is a scalar field (with light-like gradient) which identifies the past light-cone centered on the observer and spanned by the null momentum $k_\mu=\pa_\mu V$ of the light signals emitted by the sources, and $A(x)$ is a scalar field (with a time-like gradient) associated with the following unit vector:
\beq
n_\mu=\frac{\pa_\mu A}{|\pa_\alpha A\pa^\alpha A|^{1/2}},
\label{2}
\eeq
which possibly represents a convenient four-velocity reference,
{{but} %MDPI: Please confirm if the bold is unnecessary and can be removed. The following highlights are the same.
 in general depending on the particular observations we are interested in.}
The particular choice of $\rho$, $A$, and $B$ obviously depends on the physical situation and on the type of observation being performed.

Suppose now that we are interested in sources localized between the constant redshift surfaces $z=z_s$ and $z=z_s+\Delta z_s$. In this case, we can choose $B=1+z$, where $z$ is the standard redshift parameter defined in general by
\beq
1+z=\frac{\left( u_\mu k^\mu \right)_s}{\left( v_\mu k^\mu \right)_o}\,,
\label{3}
\eeq
where $u_\mu$ and $v_\mu$ are the velocities of the source and the observer, respectively, {{not} necessarily co-moving in the given geometry}, and the subscripts ``$s$'' and ``$o$'' respectively denote the source and observer positions (see, e.g., \cite{2}). In this case, as we show below, we find that our integral (\ref{1}) can exactly reproduce the standard number-count prescription of \cite{2,3,4}, provided the fields $A$ and $\rho$ are appropriately chosen, {{as} is shown explicitly just after Equation (\ref{14})}.

It is convenient for this purpose to work in the so-called Geodesic Light-Cone (GLC) gauge based on the coordinates $x^\mu=\left( \tau,w,\theta^a \right)$ and $a=1,2$, where the most general cosmological metric is parameterized by six arbitrary functions $\U$, $U_a$, $\gamma_{ab}=\gamma_{ba}$, and the line-element takes the following form \cite{6}:
\beq
ds^2_{GLC}=-2\U dwd\tau+\U^2dw^2+\ga_{ab}\left(d\theta^a-U^a dw\right)\left(d\theta^b-U^b dw\right).
\label{4}
\eeq
{Let} us recall here for the reader's convenience that $w$ is a null coordinate ($\pa_\mu w \pa^\mu w=0$), that in this gauge the light signals travel along geodesics with constant $w$ and $\theta^a$, and that the time coordinate $\tau$ coincides with the time of the synchronous gauge \cite{7}. In fact, we can easily determine $\pa_\mu \tau$ defines a geodesic flow, i.e., that $(\pa^\nu \tau) \nabla_\nu (\pa_\mu \tau)=0$, which is in agreement with the condition $g^{\tau\tau}=-1$ following from the metric (\ref{4}). 

Working in the GLC gauge, we can identify $V=w$; thus, $k_\mu=\pa_\mu w$ and $k^\mu=-\U^{-1}\delta^\mu_\tau$. Moreover, we can conveniently impose the {\em {general} %MDPI: Please confirm if the italics is unnecessary and can be removed. The following highlights are the same.
 temporal gauge} defined by the condition\endnote{This choice generalises the definition of the temporal gauge, already introduced in \cite{8}, such that it can be applied to the case of an arbitrary observer velocity $v_\mu$.}
\beq
\left( v_\mu k^\mu \right)_o=-1\,,
\label{5}
\eeq
where the past light cones $w=w_o=${const} are simply labeled by the reception time $\tau_o$ of the corresponding light signals \cite{8}, i.e., $w_o=\tau_o$. Hence, in this gauge we have
\beq
1+z=\left( u_\tau \U^{-1} \right)_s\
\label{6}
\eeq
and we can replace the $\tau$ integration of Equation~(\ref{1}) with the $z$ integration defined by
\beq
d\tau=dz\left( \frac{d\tau}{dz} \right)=\frac{dz}{\pa_\tau\left( u_\tau \U^{-1} \right)}\,.
\label{7}
\eeq
{Note} that Equation~\eqref{7} has to be further integrated, as prescribed by Equation~\eqref{1}, on the spatial hypersurface containing the averaged source. 
Thus, in order to avoid any confusion between the integrated quantities and the boundaries of the integrals, we omit the explicit subscript ``$s$'' in the differential integration measure. Finally, for the metric (\ref{4}) we have $\sqrt{-g}=\U\sqrt{\gamma}$, where $\gamma=\det\gamma_{ab}$. Thus, our integral prescription (\ref{1}) takes the form
\bea
I(\Delta z_s)&=&-\int d\tau\,dw\,d^2\theta\,\sqrt{\gamma}\,\rho\,\delta\left( w_o-w \right)\,
\Theta(z_s+\Delta z_s-z)\,\Theta(z- z_s)\,n_\tau\nonumber\\
&=&-\int dz\,d^2\theta\,\sqrt{\gamma}\,\rho\,
\Theta(z_s+\Delta z_s-z)\,\Theta(z- z_s)\,\frac{n_\tau}{\pa_\tau\left( u_\tau \U^{-1} \right)}\nonumber\\
&=&-\int_{z_s}^{z_s+\Delta z_s} dz\,d^2\theta\,\sqrt{\gamma}\,\rho
\,\frac{n_\tau}{\pa_\tau\left( u_\tau \U^{-1} \right)}\,,
\label{8}
\eea
where we have used the explicit form of $n_\mu$ of Equation~(\ref{2}). It should be stressed here that $u_\tau$ is the time component of the source velocity, while for the moment $n_\tau$ and $\rho$ are both arbitrary variables to be adapted to the physical situation under consideration.\endnote{In our previous paper \cite{1}, we applied the above integral in the limit of the small redshift bin $\Da z_s \ra 0$.} Finally, all the integrated functions have to be evaluated on the past light cone $w=w_o$.

We now Consider the number count integral, which evaluates the number of sources $dN$ located inside an infinitesimal layer of thickness $d\lambda$ at a distance $d_A(z)$ and seen by a given observer within a bundle of null geodesics subtending the solid angle $d\Omega$ \cite{2,3}:
\beq
dN=n\,dV\equiv n\,d\Omega\,d\lambda\,d^2_A\,\left( -u_\mu k^\mu \right)\,.
\label{9}
\eeq
{Here,} $n$ is the number density of the sources per unit proper volume, $d_A$ is the angular distance, and $u_\mu$ (as before) is the velocity field of the given sources. Finally, $\lambda$ is a scalar affine parameter along the path $x^\mu\left( \lambda \right)$ of the light signals such that $k^\mu=d x^\mu/d\lambda$, and is normalized along the observer world-line by the condition 
\beq
\left( v_\mu k^\mu \right)_o=-1\,,
\label{10}
\eeq
where $v_\mu$ is the observer velocity and the scalar product is evaluated at the observer position \cite{3}.

We now move to the GLC coordinates, where $k^\mu=g^{\mu\nu}\pa_\nu w=-\U^{-1}\delta^\mu_\tau$. The condition $k^\mu=dx^\mu/d\lambda$ then provides the {following:} %MDPI: Please confirm that the space before the comma is necessary and that the two commas (,  \qquad,) are correct
\beq
\frac{d\tau}{d\lambda}=-\U^{-1},\qquad \qquad\frac{dw}{d\lambda}=0,  
\qquad \qquad \frac{d\theta^a}{d\lambda}=0\,,
\label{11}
\eeq
and the normalization condition (\ref{10}) exactly coincides with the general temporal gauge (\ref{5}). In these coordinates, as anticipated in \cite{9,10}, the angular distance $d_A$ satisfies 
\beq
d^2_A\,d\Omega=\sqrt{\gamma}\,d^2\ti\theta\,;
\label{12}
\eeq
see Appendix \ref{appa} for an explicit derivation of the above equation.
Here, $\ti \theta^a$ represents the angular coordinates of the so-called ``observational gauge'' \cite{11,12}, defined by exploiting the residual gauge freedom of the GLC gauge in such a way that the angular directions exactly coincide with those expressed in a system of Fermi Normal Coordinates (FNC).\endnote{The angular directions related to local observations (also used in \cite{2,3}) are indeed those measured by a free-falling observer, and can be identified with the angles of the FNC system \cite{11} where the metric is locally flat around all points of a given world line, with leading curvature corrections (which are quadratic) in the distance.} Using
\beq
d\lambda=dz\left( \frac{dz}{d\tau} \right)^{-1}\left( \frac{d\tau}{d\lambda} \right)^{-1}\,
\label{13}
\eeq
and applying Equations (\ref{7}) and (\ref{11}), we find that the number of sources of Equation~(\ref{9}) integrated between $z_s$ and $z_s+\Delta z_s$ (as was the case before; see Equation~(\ref{8})) finally reduces to
\beq
N=-\int_{z_s}^{z_s+\Delta z_s}dz\,d^2\ti\theta\,\sqrt{\gamma}\,n\frac{u_\tau}{\pa_\tau\left( u_\tau\U^{-1} \right)}\,.
\label{14}
\eeq

We are now in the position of comparing this result with our previous average integral~(\ref{8}). It is clear that the two are exactly the same, provided the following three conditions are satisfied: ({$i$}%MDPI: Please confirm if the italics is unnecessary and can be removed. The following highlights are the same
) our scalar density $\rho$ coincides with the number density $n$; ({$ii$}) the scalar parameter $A$ is chosen in such a way that the unit vector $n_\mu$ coincides with the source velocity $u_\mu$ (and, obviously, $n_\tau= u_\tau$); and ({$iii$}) the angular directions are expressed in terms of the angles fixed by the observational gauge, i.e., $\theta^a \ra \ti\theta^a$.

It may be appropriate to recall at this point that the number count prescription has been recently used in (apparently) different forms by other authors. For instance, in the context of defining physically appropriate averaging prescriptions the number count has been presented in the following form \cite{4}:
\beq
dN=n\,dV={n\, d^2_A\over (1+z)\,H_{||}} \,dz\,d\Om\,,
\label{15}
\eeq
where $H_{||}$ is a local longitudinal expansion parameter defined by
\beq
H_{||}\equiv\left( 1+z \right)^{-2}\,k^\mu k^\nu\nabla_\mu u_\nu\,.
\label{16}
\eeq
{Moving} to the GLC coordinates and using the temporal gauge (\ref{5}), we can easily obtain
\beq
(1+z)\,H_{||}=-u^{-1}_\tau\,\pa_\tau\left( u_\tau\U^{-1} \right)\,.
\label{17}
\eeq
{By }inserting this result in the definition in (\ref{15}) and using Equation~(\ref{11})  in Equation~(\ref{9}), we can immediately determine that the 
number count expressions in (\ref{15}) and (\ref{9}) are exactly the same.

As a second example, recall the form of the number-count integral presented in~\cite{5} based on the following volume element:
\beq
dV=\sqrt{-g}\,\epsilon_{\mu\nu\alpha\beta}\,u^\mu dx^\nu dx^\alpha dx^\beta\,,
\label{18}
\eeq
where, as before, $u^\mu$ is the source velocity. Moving to the GLC gauge and projecting this volume element on the light cone $w=\rm{const}$, $dw=0$, we obtain
\beq
dV=\U\,\sqrt{\gamma}\,u^wd\tau d^2\theta\,.
\label{19}
\eeq
{Recalling} that $u^w=g^{w\nu}u_\nu=-u_\tau\U^{-1}$ and again using the expression $d\tau/dz$ provided in Equation~(\ref{7}), it is immediately clear that Equation~(\ref{19}) reduces to the same expression of the number count integrand in (\ref{14}), provided we add the source density $n$ and, as before, the GLC angles are identified with those of the observational gauge \cite{11,12} $d^2\theta \ra d^2 \ti\theta$. 

The discussion presented in this paper provides a further example of the crucial role played by the GLC coordinates in the simplification and comparison of formal non-perturbative expressions of physical observables. In addition, the simple expression obtained here for the volume element $dV$ in terms of observable variables such as the redshift and observation angles is promising for a number of different physical applications, which will be discussed in forthcoming papers.

We would finally remark that in this brief article we have expressed the galaxy number counts in terms of the redshift. A recent interesting paper \cite{Fonseca:2023uay} has proposed studying galaxy number counts as a function of the luminosity distance of the given sources, rather than their redshift, and showed that there are already differences between the two computational methods at the first perturbative order. Thus, in the near future we plan to evaluate the exact expression of the galaxy number counts in terms of the luminosity distance by applying the covariant averaging formalism and using the GLC coordinate approach presented in this~paper.

%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\vspace{6pt} 

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\authorcontributions{{Conceptualization, G. Fanizza, M. Gasperini and G. Marozzi; methodology,  G. Fanizza, M. Gasperini and G. Marozzi; formal analysis,  G. Fanizza, M. Gasperini and G. Marozzi; original draft preparation,  G. Fanizza, M. Gasperini and G. Marozzi; review and editing,  G. Fanizza, M. Gasperini and G. Marozzi. 
All authors} have read and agreed to the published version of the manuscript.}
%MDPI: For research articles with several authors, a short paragraph specifying their individual contribu-tions must be provided. The following statements should be used “Conceptualization, X.X. and Y.Y.; methodology, X.X.; software, X.X.; validation, X.X., Y.Y. and Z.Z.; formal analysis, X.X.; investiga-tion, X.X.; resources, X.X.; data curation, X.X.; writing—original draft preparation, X.X.; writ-ing—review and editing, X.X.; visualization, X.X.; supervision, X.X.; project administration, X.X.; funding acquisition, Y.Y. All authors have read and agreed to the published version of the manu-script.” Please turn to the CRediT taxonomy for the term explanation. Authorship must be limited to those who have contributed substantially to the work reported.

\funding{This research received no external funding.}

\institutionalreview{Not applicable.}

\informedconsent{Not applicable.}

\dataavailability{Not applicable.} 

\acknowledgments{{G.} %MDPI: Please ensure that all individuals included in this section have consented to the acknowledgement.
 Fanizza acknowledges support by Funda\c{c}\~{a}o para a Ci\^{e}ncia e a Tecnologia (FCT) under the program {\it {``Stimulus"} %MDPI: Please confirm if the italics is unnecessary and can be removed. The following highlights are the same
} with the grant no. CEECIND/04399/2017/CP1387/CT0026, and through the research project with ref. number PTDC/FIS-AST/0054/2021.
M. Gasperini and G. Marozzi are supported in part by INFN under the program TAsP ({\it {``Theoretical Astroparticle Physics"}}). M. Gasperini is also supported by the research grant number 2017W4HA7S {\it {``NAT-NET: Neutrino and Astroparticle Theory Network"}}, under the program PRIN 2017 funded by the Italian Ministero dell'Universit\`a e della Ricerca (MUR). G. Fanizza and M. Gasperini wish to thank the kind hospitality and support of the TH Department of CERN, where part of this work has been carried out. Finally, we are very grateful to Gabriele Veneziano for his fundamental contribution and collaboration during the early stages of this work.}

\conflictsofinterest{The authors declare no conflict of interest.} 

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%% Optional
%\sampleavailability{Samples of the compounds ... are available from the authors.}

\abbreviations{Abbreviations}{
The following abbreviations are used in this manuscript:\\

\noindent 
\begin{tabular}{@{}ll}
%\hl{MDPI} %MDPI: the abbreviation ``MDPI'' is not required. please confirm if it can be removed
 %& Multidisciplinary Digital Publishing Institute\\
%DOAJ & Directory of open access journals\\
%TLA & Three letter acronym\\
%LD & Linear dichroism
GLC & Geodesic Light-Cone\\
FNC & Fermi Normal Coordinates
\end{tabular}
}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%% Optional
\appendixtitles{no} % Leave argument "no" if all appendix headings stay EMPTY (then no dot is printed after "Appendix A"). If the appendix sections contain a heading then change the argument to "yes".
\appendixstart
\appendix
\section[\appendixname~\thesection]{}\label{appa}
%\subsection[\appendixname~\thesubsection]{}

In order to prove Equation~(\ref{12}), we can combine Equations (3.15) and (3.17) from~\cite{10} to express the angular distance in generic GLC coordinates as follows:
\beq
d_A^2 = 4 v_\tau^2 \sqrt{ \gamma_s} \left[\frac{{\rm det} \,(\partial_\tau \gamma_{ab})}{\sqrt{\gamma}} \right]_o^{-1}
\eeq
where, as before, the subscripts ``$s$'' and ``$o$'' respectively denote the source and observer positions. 
We can now move to the observational gauge \cite{11},  
  where we choose a system at rest with the local observer ($v_\tau^2=1$) and use Equation (3.16) from~\cite{11} to obtain, after a simple calculation of $\gamma_{ab}$, 
 \beq
 \ti{d}_A^2 =  \frac{\sqrt{\gamma_s}}{\sin \ti{\theta}}\, ,
 \eeq
from which we have
 \beq
 \ti{d}_A^2  \,d \ti{\Omega} =  \sqrt{\gamma_s} \,d^2  \ti{\theta},
 \eeq
where the tilde denotes the variables of the observational gauges. 
Finally, note that by using Equations~(3.13)--(3.15) from
\cite{10} it can easily be shown that the left-hand side of the above equation does not depend on the particular choice of $v_\tau$.

%\section[\appendixname~\thesection]{}
%All appendix sections must be cited in the main text. In the appendices, Figures, Tables, etc. should be labeled, starting with ``A''---e.g., Figure A1, Figure A2, etc.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{adjustwidth}{-\extralength}{0cm}
\printendnotes[custom] % Un-comment to print a list of endnotes

\reftitle{References}

\PublishersNote{}
\end{adjustwidth}
\end{document}